# On the conjugates of Christoffel words

### v4

@@ 8.2. Consequences
If $b_{m-1}-d_{m-1}\geq 1$ and $d_{m-2}=0$, then by Lemmas 8.14 and 8.15, $B=V_{m-2}^{h+1}$ if $b_{m-1}-d_{m-1}<d_{m-1}+1$ and $V_{m-2}$ is a prefix of $V_{m-3}V_{m-2}$, and $B=V_{m-2}^{h}$ otherwise. But, by Corollary 8.10 with $m$ replaced by m1m-1, $V_{m-2}$ is a prefix of $V_{m-3}V_{m-2}$ if and only if the sequence $d_{1},\ldots,d_{m-2}$ is not alternating.
If $b_{m-1}-d_{m-1}\geq 1$ and $d_{m-2}=b_{m-2}$, then by Lemma 8.18 with $m$ replaced by m2m-2, $V_{m-2}$ is not a suffix of $V_{m-2}V_{m-3}$. Thus by Lemmas 8.14 and 8.15, $B=V_{m-2}^{h}$.
$(vii)$ We have $m\geq 3$: indeed, for $m=2$, $V_{2}=ba^{b_{1}+1}$ is a Christoffel word, which was excluded. Since $d_{m-1}=b_{m-1}$, then $d_{m-2}=0$ by the greedy condition, and $V_{m}=V_{m-3}V_{m-2}^{b_{m-1}+1}$. If $V_{m-2}$ is not a prefix of $V_{m-3}V_{m-2}$, then by Corollary 8.10, the sequence $d_{1},\ldots,d_{m-2}$ is alternating; then, since $d_{m-2}=0,d_{m-1}=b_{m-1},d_{m}=0$, the sequence $d_{1},\ldots,d_{m}$ is alternating too, and $V_{m}$ is a Christoffel word, a contradiction. Thus $V_{m-2}$ is a prefix of $V_{m-3}V_{m-2}$, and by Lemmas 8.13 and 8.16, we get that $B=V_{m-2}$. ∎
9. The Sturmian graph revisited
@@
The last assertion follows from the equality $q_{m}-2=\sum_{1\leq i\leq m}c_{i}q_{i-1}$, see Lemma 10.2 (iii). ∎
The previous corollary has a graph-theoretic interpretation. We construct an edge-labelled directed graph $(V,E)$, that we shall call compact graph for short. It will turn out to be a graph introduced in [EMSV], where it is called the compact directed acyclic word graph of $p$.
For the construction of this graph, it is convenient to view (9.1) as a word over the letters $L_{0},\ldots,L_{m-1}$; in particular we consider prefixes of this word, which are the elements of $V$; the latter set has therefore $c_{1}+\cdots+c_{m}+1$ elements. We denote by $1$ the vertex corresponding to the empty word. For each vertex $UL_{i},0\leq i\leq m-1$, there is an edge labelled $L_{i}$ from $U$ to $UL_{i}$:
@@
Figure 1. The compact graph $(V,E)$
@@
Let now $s$ be any suffix of $p$. Then by Corollary 9.1, $s$ is equal to $L_{0}^{d_{1}}L_{1}^{d_{2}}\cdots L_{m-1}^{d_{m}}$, where $\sum_{1\leq i\leq m}d_{i}q_{i-1}$ is the lazy Ostrowski representation of the length of $s$. Let $k$ be maximal such that $d_{k}\neq 0$. Then $s=L_{0}^{d_{1}}L_{1}^{d_{2}}\cdots L_{k-1}^{d_{k}}$. We claim that for each $j=1,\ldots,k$, there is a path labelled $L_{0}^{d_{1}}\cdots L_{j-1}^{d_{j}}$ from the origin until the vertex $L_{0}^{c_{1}}\cdots L_{j-2}^{c_{j-1}}L_{j-1}^{d_{j}}$. The claim is clear for $j=1$, since one has the edges $1\to L_{0}\to L_{0}^{2}\to\cdots\to L_{0}^{c_{1}}$, all labelled $L_{0}$ and since $d_{1}\leq b_{1}=c_{1}$. Admitting the claim for $j\leq k-1$, we prove it for $j+1\leq k$. If $d_{j+1}=0$, then $j+1<k$ and by laziness, $d_{j}=b_{j}=c_{j}$ (the last equality holds since $j<m$); then the path for $j+1$ is the same as that for $j$: there is a path labelled $L_{0}^{d_{1}}\cdots L_{j-1}^{d_{j}}=L_{0}^{d_{1}}\cdots L_{j-1}^{d_{j}}L_{j}^{d_{j+1}}$ from the origin until the vertex $L_{0}^{c_{1}}\cdots L_{j-2}^{c_{j-1}}L_{j-1}^{d_{j}}=L_{0}^{c_{1}}\cdots L_{j-2}^{c_{j-1}}L_{j-1}^{c_{j}}L_{j}^{d_{j+1}}$. Suppose now that $d_{j+1}\neq 0$; in the graph we have the $c_{j+1}$ consecutive edges $L_{0}^{c_{1}}\cdots L_{j-2}^{c_{j-1}}L_{j-1}^{d_{j}}\to L_{0}^{c_{1}}\cdots L_{j-2}^{c_{j-1}}L_{j-1}^{c_{j}}L_{j}\to L_{0}^{c_{1}}\cdots L_{j-2}^{c_{j-1}}L_{j-1}^{c_{j}}L_{j}^{2}\to\cdots\to L_{0}^{c_{1}}\cdots L_{j-2}^{c_{j-1}}L_{j-1}^{c_{j}}L_{j}^{c_{j+1}}$, all labelled $L_{j}$; note that $d_{j+1}\leq c_{j+1}$: indeed, the representation is legal, hence $d_{j+1}\leq b_{j+1}$ and $b_{j+1}=c_{j+1}$, except if $j+1=m$; but in this case, since $s$ is of length at most $q_{m}-2$, we have $d_{m}\leq b_{m}-1=c_{m}$ by Corollary 10.3; thus the claim follows for $j+1$ too.
Thus, for $j=k$, we obtain that there is a path starting from the origin and labelled $s$, in the graph. ∎
@@
The compact graph is the Sturmian graph of [EMSV, EFGMS]. This will be verified now.
@@
In the light of [EMSV] (Theorem 19, and beginning of Section 19, where compaction is described [^2] The notion of compaction of an automaton appears in [BBHME].), this proves that the graph of Corollary 9.3 is the Sturmian graph. It implies also Theorem 47 of [EFGMS]: each path in the Sturmian graph, with label $N$, corresponds to the lazy Ostrowski representation of $N$.
We indicate now how to construct the compact graph using the iterated palindromization of Aldo de Luca [dL] (see also [Re, Section 12.1]). Recall the definition of this operator, denoted $Pal$. One defines first the right palindromic closure of a word $w$, denoted $w^{(+)}$: it is the shortest palindrome having $w$ as prefix. Then the mapping $Pal$ from a free monoid into itself is defined recursively by $Pal(1)=1$ and $Pal(wx)=(Pal(w)x)^{(+)}$ for any word $w$ and any letter $x$. The theorem of de Luca is that $Pal$ is a bijection from $\{a,b\}^{*}$ onto the set of central words.
@@
The central word $p$ has the directive word $v=a^{c_{1}}b^{c_{2}}a^{c_{3}}\cdots(a\,\mbox{or}\,\,b)^{c_{m}}$. The word $p=Pal(v)$ has $1+c_{1}+\cdots+c_{m}$ palindromic prefixes, which are the formal prefixes of (9.1). In particular, $L_{i}=Pal(a^{c_{1}}\cdots(a\,\mbox{or}\,\,b)^{c_{i}})^{-1}Pal(a^{c_{1}}\cdots(a\,\mbox{or}\,\,b)^{c_{i}}(b\,\mbox{or}\,\,a))$.
Proof.
We know that the Slope of $M_{m}$, and in particular of the lower Christoffel word in the conjugation class of $M_{m}$, is $S=[0,a_{1},\ldots,a_{n}]$ (Theorem 7.3). It follows from the analysis at the end of Section 6 that $s=[0,b_{1},\ldots,b_{m}]$ if $b_{1}\geq 1$, and $s=[b_{2},\ldots,b_{m}]$ if $b_{1}=0$. It follows from Theorem 14.2.3 in [Re] (the result is from [GKP, Section 14.2.3]) that the path leading from the root to the node $s$ in the Stern-Brocot tree is coded by the word $v=a^{c_{1}}b^{c_{2}}\cdots(a\,\mbox{or}\,\,b)^{c_{m}}$ ($a$ means left, and $b$ means right). It follows then from the correspondence between the Stern-Brocot tree, the tree of Christoffel words, and the tree of central words (see Sections 12.1, 14.1 and 14.2 in [Re]) that the path from the root to $p$ in the latter tree is coded by $v$, proving the first assertion.
The word $p=Pal(v)$ has $|v|+1$ palindromic prefixes. By Lemma 7.7, all the words $M_{i-1}^{c}M_{i-2}^{c_{i-1}}\cdots M_{1}^{c_{2}}M_{0}^{c_{1}}$, where $i=1,\ldots,m$, $0\leq c\leq c_{i}$, are palindromes. Hence their reversals $L_{0}^{c_{1}}L_{1}^{c_{2}}\ldots L_{i-2}^{c_{i-1}}L_{i-1}^{c}$ are palindromes too, and are suffixes of $p$ by Corollary 9.1, proving the second assertion.
The last assertion then follows. ∎
@@
Proof.
Consider some node on the tree of central words; as in the proof above, we associate with it the word $v\in\{a,b\}^{*}$, which encodes the path from the root to this vertex; then this vertex is $Pal(v)$ (see [Re, Section 12.1]). The construction of the compact graph then follows from the proposition. And from this the construction of the Sturmian graph also follows. ∎
@@
Figure 2. A path in the tree of central words, a compact graph and a Sturmian graph
10. Appendix: a proof of existence and uniqueness of the greedy and lazy representations
@@
(10.1) $N=d_{1}q_{0}+d_{2}q_{1}+\cdots+d_{k}q_{k-1}.$
(i) If in (10.1) the sequence $d_{i}$ is alternating, with $k=0$ or $d_{k}\neq 0$, then $N=q_{k}-1$.
(ii) If in (10.1) one assumes that the representation is lazy and that $k=0$ or $d_{k}=b_{k}$, then $N\geq q_{k}-1$.
Note that we say that the representation (10.1) is legal (resp. greedy, resp. lazy) if the representation (3.2) of $N$ obtained by letting $d_{i}=0$ for $i=k+1,\ldots,m$ has this property.
Proof.
@@
(ii) This is clearly true for $k=0$ and $k=1$, since $q_{0}=1$ and $b_{1}=a_{1}-1=q_{1}-1$. Assume that $k\geq 1$ and that it is true for k1k-1 and $k$, and we prove it for $k+1$; thus we consider a sequence $d_{1},\ldots,d_{k+1}$ with $d_{k+1}=b_{k+1}$. If $d_{k}\neq 0$, then $N=\sum_{i=1}^{k+1}d_{i}q_{i-1}\geq b_{k+1}q_{k}+q_{k-1}=a_{k+1}q_{k}+q_{k-1}=q_{k+1}\geq q_{k+1}-1$. If $d_{k}=0$ then, assuming that $k\geq 2$, we have $d_{k-1}=b_{k-1}$ since the representation is lazy ; thus by induction, $N\geq b_{k+1}q_{k}+q_{k-1}-1=a_{k+1}q_{k}+q_{k-1}-1=q_{k+1}-1$. The remaining case is $k=1,d_{2}=b_{2}=a_{2},d_{1}=0$ and $N=d_{2}q_{1}=a_{2}a_{1}=q_{2}-1$. ∎
@@
Proof.
(i) We prove the first inequality by induction on $k$. For $k=0$, $N=0$ and it holds since $q_{0}=1$. For $k=1$, it holds since $d_{1}\leq a_{1}-1$. Assume that $k\geq 1$, and that it is true for $1,\ldots,k$, and we prove it for $k+1$. If $d_{k+1}=a_{k+1}=b_{k+1}$ (since $k+1\geq 2$), then $d_{k}=0$ by (3.4); then $d_{1}q_{0}+d_{2}q_{1}+\cdots+d_{k+1}q_{k}=d_{1}q_{0}+d_{2}q_{1}+\ldots+d_{k-1}q_{k-2}+a_{k+1}q_{k}\leq$ (by induction) $q_{k-1}-1+a_{k+1}q_{k}=q_{k+1}-1$; if on the other hand, $d_{k+1}\leq a_{k+1}-1$, then $d_{1}q_{0}+d_{2}q_{1}+\cdots+d_{k+1}q_{k}=d_{1}q_{0}+d_{2}q_{1}+\cdots+d_{k}q_{k-1}+d_{k+1}q_{k}\leq$ (by induction) $q_{k}-1+a_{k+1}q_{k}-q_{k}<-1+q_{k-1}+a_{k+1}q_{k}=q_{k+1}-1$.
@@
(iii) is proved similarly by induction. ∎
@@
Suppose now that $d_{k}\neq b_{k}$; then $d_{k}\leq b_{k}-1$. Then $N+q_{k-1}$ has the lazy representation $(d_{k}+1)q_{k-1}+\sum_{i=1}^{k-1}d_{i}q_{i-1}$. Thus by Lemma 10.2 (ii), $N+q_{k-1}\leq q_{k}+q_{k-1}-2$, hence $N\leq q_{k}-2$. ∎
Proof of Proposition 3.1.
@@
For the claim, we may assume that $k\geq 2$; we have $N^{\prime}=\sum_{i=1}^{k}(b_{i}-d_{i})q_{i-1}$ by Lemma 10.2 (iii). We have $N^{\prime}=(b_{k}-d_{k})q_{k-1}+r$, where $r=\sum_{i=1}^{k-1}(b_{i}-d_{i})q_{i-1}$. By lazyness, this is a greedy representation of $r$. Hence $r\leq q_{k-1}-1$ by Lemma 10.2 (i); since $r\geq 0$, the claim follows. ∎
Acknowledgments We thank the two anonymous referees for their comments. This work was partially supported by NSERC, Canada.

### v5

@@ 8.2 Consequences
If $b_{m-1}-d_{m-1}\geq 1$ and $d_{m-2}=0$, then by Lemmas 8.25 and 8.27, $B=V_{m-2}^{h+1}$ if $b_{m-1}-d_{m-1}<d_{m-1}+1$ and $V_{m-2}$ is a prefix of $V_{m-3}V_{m-2}$, and $B=V_{m-2}^{h}$ otherwise. But, by Corollary 8.17 with $m$ replaced by m1m-1, $V_{m-2}$ is a prefix of $V_{m-3}V_{m-2}$ if and only if the sequence $d_{1},\ldots,d_{m-2}$ is not alternating.
If $b_{m-1}-d_{m-1}\geq 1$ and $d_{m-2}=b_{m-2}$, then by Lemma 8.33 with $m$ replaced by m2m-2, $V_{m-2}$ is not a suffix of $V_{m-2}V_{m-3}$. Thus by Lemmas 8.25 and 8.27, $B=V_{m-2}^{h}$.
$(vii)$ We have $m\geq 3$: indeed, for $m=2$, $V_{2}=ba^{b_{1}+1}$ is a Christoffel word, which was excluded. Since $d_{m-1}=b_{m-1}$, then $d_{m-2}=0$ by the greedy condition, and $V_{m}=V_{m-3}V_{m-2}^{b_{m-1}+1}$. If $V_{m-2}$ is not a prefix of $V_{m-3}V_{m-2}$, then by Corollary 8.17, the sequence $d_{1},\ldots,d_{m-2}$ is alternating; then, since $d_{m-2}=0,d_{m-1}=b_{m-1},d_{m}=0$, the sequence $d_{1},\ldots,d_{m}$ is alternating too, and $V_{m}$ is a Christoffel word, a contradiction. Thus $V_{m-2}$ is a prefix of $V_{m-3}V_{m-2}$, and by Lemmas 8.23 and 8.29, we get that $B=V_{m-2}$.
9 The Sturmian graph revisited
@@
The last assertion follows from the equality $q_{m}-2=\sum_{1\leq i\leq m}c_{i}q_{i-1}$, see Lemma 10.3 (iii).
The previous corollary has a graph-theoretic interpretation. We construct an edge-labelled directed graph $(V,E)$, that we shall call compact graph for short. It will turn out to be a graph introduced in Epifanio et al. (2007), where it is called the compact directed acyclic word graph of $p$.
For the construction of this graph, it is convenient to view (17) as a word over the letters $L_{0},\ldots,L_{m-1}$; in particular we consider prefixes of this word, which are the elements of $V$; the latter set has therefore $c_{1}+\cdots+c_{m}+1$ elements. We denote by $1$ the vertex corresponding to the empty word. For each vertex $UL_{i},0\leq i\leq m-1$, there is an edge labelled $L_{i}$ from $U$ to $UL_{i}$:
@@
Figure 1: The compact graph $(V,E)$
@@
Let now $s$ be any suffix of $p$. Then by Corollary 9.1, $s$ is equal to $L_{0}^{d_{1}}L_{1}^{d_{2}}\cdots L_{m-1}^{d_{m}}$, where $\sum_{i=1}^{m}d_{i}q_{i-1}$ is the lazy Ostrowski representation of the length of $s$. Let $k$ be maximal such that $d_{k}\neq 0$. Then $s=L_{0}^{d_{1}}L_{1}^{d_{2}}\cdots L_{k-1}^{d_{k}}$. We claim that for each $j=1,\ldots,k$, there is a path labelled $L_{0}^{d_{1}}\cdots L_{j-1}^{d_{j}}$ from the origin until the vertex $L_{0}^{c_{1}}\cdots L_{j-2}^{c_{j-1}}L_{j-1}^{d_{j}}$. The claim is clear for $j=1$, since one has the edges $1\to L_{0}\to L_{0}^{2}\to\cdots\to L_{0}^{c_{1}}$, all labelled $L_{0}$ and since $d_{1}\leq b_{1}=c_{1}$. Admitting the claim for $j\leq k-1$, we prove it for $j+1\leq k$. If $d_{j+1}=0$, then $j+1<k$ and by laziness, $d_{j}=b_{j}=c_{j}$ (the last equality holds since $j<m$); then the path for $j+1$ is the same as that for $j$: there is a path labelled $L_{0}^{d_{1}}\cdots L_{j-1}^{d_{j}}=L_{0}^{d_{1}}\cdots L_{j-1}^{d_{j}}L_{j}^{d_{j+1}}$ from the origin until the vertex $L_{0}^{c_{1}}\cdots L_{j-2}^{c_{j-1}}L_{j-1}^{d_{j}}=L_{0}^{c_{1}}\cdots L_{j-2}^{c_{j-1}}L_{j-1}^{c_{j}}L_{j}^{d_{j+1}}$. Suppose now that $d_{j+1}\neq 0$; in the graph we have the $c_{j+1}$ consecutive edges $L_{0}^{c_{1}}\cdots L_{j-2}^{c_{j-1}}L_{j-1}^{d_{j}}\to L_{0}^{c_{1}}\cdots L_{j-2}^{c_{j-1}}L_{j-1}^{c_{j}}L_{j}\to L_{0}^{c_{1}}\cdots L_{j-2}^{c_{j-1}}L_{j-1}^{c_{j}}L_{j}^{2}\to\cdots\to L_{0}^{c_{1}}\cdots L_{j-2}^{c_{j-1}}L_{j-1}^{c_{j}}L_{j}^{c_{j+1}}$, all labelled $L_{j}$; note that $d_{j+1}\leq c_{j+1}$: indeed, the representation is legal, hence $d_{j+1}\leq b_{j+1}$ and $b_{j+1}=c_{j+1}$, except if $j+1=m$; but in this case, since $s$ is of length at most $q_{m}-2$, we have $d_{m}\leq b_{m}-1=c_{m}$ by Corollary 10.5; thus the claim follows for $j+1$ too.
Thus, for $j=k$, we obtain that there is a path starting from the origin and labelled $s$, in the graph.
@@
The compact graph is the Sturmian graph of Epifanio et al. (2007, 2012). This will be verified now.
@@
In the light of Epifanio et al. (2007) (Theorem 19, and beginning of Section 19, where compaction is described [^2] The notion of compaction of an automaton appears in Blumer et al. (1987).), this proves that the graph of Corollary 9.5 is the Sturmian graph. It implies also Theorem 47 of Epifanio et al. (2012): each path in the Sturmian graph, with label $N$, corresponds to the lazy Ostrowski representation of $N$.
We indicate now how to construct the compact graph using the iterated palindromization of Aldo de Luca de Luca (1997) (see also (Reutenauer, 2019, Section 12.1)). Recall the definition of this operator, denoted $Pal$. One defines first the right palindromic closure of a word $w$, denoted $w^{(+)}$: it is the shortest palindrome having $w$ as prefix. Then the mapping $Pal$ from a free monoid into itself is defined recursively by $Pal(1)=1$ and $Pal(wx)=(Pal(w)x)^{(+)}$ for any word $w$ and any letter $x$. The theorem of de Luca is that $Pal$ is a bijection from $\{a,b\}^{*}$ onto the set of central words.
@@
The central word $p$ has the directive word $v=a^{c_{1}}b^{c_{2}}a^{c_{3}}\cdots(a\,\mbox{or}\,\,b)^{c_{m}}$. The word $p=Pal(v)$ has $1+c_{1}+\cdots+c_{m}$ palindromic prefixes, which are the formal prefixes of (17). In particular, $L_{i}=Pal(a^{c_{1}}\cdots(a\,\mbox{or}\,\,b)^{c_{i}})^{-1}Pal(a^{c_{1}}\cdots(a\,\mbox{or}\,\,b)^{c_{i}}(b\,\mbox{or}\,\,a))$.
Proof 9.7.
We know that the Slope of $M_{m}$, and in particular of the lower Christoffel word in the conjugation class of $M_{m}$, is $S=[0,a_{1},\ldots,a_{n}]$ (Theorem 7.5). It follows from the analysis at the end of Section 6 that $s=[0,b_{1},\ldots,b_{m}]$ if $b_{1}\geq 1$, and $s=[b_{2},\ldots,b_{m}]$ if $b_{1}=0$. It follows from Theorem 14.2.3 in Reutenauer (2019) (the result is from (Graham et al., 1989, Section 14.2.3)) that the path leading from the root to the node $s$ in the Stern-Brocot tree is coded by the word $v=a^{c_{1}}b^{c_{2}}\cdots(a\,\mbox{or}\,\,b)^{c_{m}}$ ($a$ means left, and $b$ means right). It follows then from the correspondence between the Stern-Brocot tree, the tree of Christoffel words, and the tree of central words (see Sections 12.1, 14.1 and 14.2 in Reutenauer (2019)) that the path from the root to $p$ in the latter tree is coded by $v$, proving the first assertion.
The word $p=Pal(v)$ has $|v|+1$ palindromic prefixes. It follows from Lemma 7.13 that all the words $M_{i-1}^{c}M_{i-2}^{c_{i-1}}\cdots M_{1}^{c_{2}}M_{0}^{c_{1}}$, where $i=1,\ldots,m$, $0\leq c\leq c_{i}$, are palindromes. Hence their reversals $L_{0}^{c_{1}}L_{1}^{c_{2}}\ldots L_{i-2}^{c_{i-1}}L_{i-1}^{c}$ are palindromes too, and are suffixes of $p$ by Corollary 9.1, proving the second assertion.
The last assertion then follows.
@@
Proof 9.9.
Consider some node on the tree of central words; as in the proof above, we associate with it the word $v\in\{a,b\}^{*}$, which encodes the path from the root to this vertex; then this vertex is $Pal(v)$ (see (Reutenauer, 2019, Section 12.1)). The construction of the compact graph then follows from the proposition. And from this the construction of the Sturmian graph also follows.
@@
Figure 2: A path in the tree of central words, a compact graph and a Sturmian graph
10 Appendix: a proof of existence and uniqueness of the greedy and lazy representations
@@
$N=d_{1}q_{0}+d_{2}q_{1}+\cdots+d_{k}q_{k-1}.$ (18)
(i) If in (18) the sequence $d_{i}$ is alternating, with $k=0$ or $d_{k}\neq 0$, then $N=q_{k}-1$.
(ii) If in (18) one assumes that the representation is lazy and that $k=0$ or $d_{k}=b_{k}$, then $N\geq q_{k}-1$.
Note that we say that the representation (18) is legal (resp. greedy, resp. lazy) if the representation (4) of $N$ obtained by letting $d_{i}=0$ for $i=k+1,\ldots,m$ has this property.
Proof 10.2.
@@
(ii) This is clearly true for $k=0$ and $k=1$, since $q_{0}=1$ and $b_{1}=a_{1}-1=q_{1}-1$. Assume that $k\geq 1$ and that it is true for k1k-1 and $k$, and we prove it for $k+1$; thus we consider a sequence $d_{1},\ldots,d_{k+1}$ with $d_{k+1}=b_{k+1}$. If $d_{k}\neq 0$, then $N=\sum_{i=1}^{k+1}d_{i}q_{i-1}\geq b_{k+1}q_{k}+q_{k-1}=a_{k+1}q_{k}+q_{k-1}=q_{k+1}\geq q_{k+1}-1$. If $d_{k}=0$ then, assuming that $k\geq 2$, we have $d_{k-1}=b_{k-1}$ since the representation is lazy ; thus by induction, $N\geq b_{k+1}q_{k}+q_{k-1}-1=a_{k+1}q_{k}+q_{k-1}-1=q_{k+1}-1$. The remaining case is $k=1,d_{2}=b_{2}=a_{2},d_{1}=0$ and $N=d_{2}q_{1}=a_{2}a_{1}=q_{2}-1$.
@@
Proof 10.4.
(i) We prove the first inequality by induction on $k$. For $k=0$, $N=0$ and it holds since $q_{0}=1$. For $k=1$, it holds since $d_{1}\leq a_{1}-1$. Assume that $k\geq 1$, and that it is true for $1,\ldots,k$, and we prove it for $k+1$. If $d_{k+1}=a_{k+1}=b_{k+1}$ (since $k+1\geq 2$), then $d_{k}=0$ by (6); then $d_{1}q_{0}+d_{2}q_{1}+\cdots+d_{k+1}q_{k}=d_{1}q_{0}+d_{2}q_{1}+\ldots+d_{k-1}q_{k-2}+a_{k+1}q_{k}\leq$ (by induction) $q_{k-1}-1+a_{k+1}q_{k}=q_{k+1}-1$; if on the other hand, $d_{k+1}\leq a_{k+1}-1$, then $d_{1}q_{0}+d_{2}q_{1}+\cdots+d_{k+1}q_{k}=d_{1}q_{0}+d_{2}q_{1}+\cdots+d_{k}q_{k-1}+d_{k+1}q_{k}\leq$ (by induction) $q_{k}-1+a_{k+1}q_{k}-q_{k}<-1+q_{k-1}+a_{k+1}q_{k}=q_{k+1}-1$.
@@
(iii) is proved similarly by induction.
@@
Suppose now that $d_{k}\neq b_{k}$; then $d_{k}\leq b_{k}-1$. Then $N+q_{k-1}$ has the lazy representation $(d_{k}+1)q_{k-1}+\sum_{i=1}^{k-1}d_{i}q_{i-1}$. Thus by Lemma 10.3 (ii), $N+q_{k-1}\leq q_{k}+q_{k-1}-2$, hence $N\leq q_{k}-2$.
Proof 10.7 (of Proposition 3.1).
@@
For the claim, we may assume that $k\geq 2$; we have $N^{\prime}=\sum_{i=1}^{k}(b_{i}-d_{i})q_{i-1}$ by Lemma 10.3 (iii). We have $N^{\prime}=(b_{k}-d_{k})q_{k-1}+r$, where $r=\sum_{i=1}^{k-1}(b_{i}-d_{i})q_{i-1}$. By lazyness, this is a greedy representation of $r$. Hence $r\leq q_{k-1}-1$ by Lemma 10.3 (i); since $r\geq 0$, the claim follows.
Acknowledgements.
We thank the two anonymous referees for their comments. This work was partially supported by NSERC, Canada.